# Video Latent Flow Matching: Optimal Polynomial Projections for Video Interpolation and Extrapolation

### v1

@@ C.3 Training Objective
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
@@
Appendix D Diffusion Transformer
In this section, we first define the Diffusion Transformer in Section D.1. Moreover, we introduce the Approximation via DiT in Section D.2.
D.1 Definitions
Definition D.1 (Multi-head self-attention).
@@ D.1 Definitions
Definition D.2 (Feed-forward).
@@
Definition D.3 (Transformer block).
Given a set of model weights $\theta^{h,m,r}=\{\{(W_{Q}^{i},W_{K}^{i},W_{V}^{i},W_{O}^{i})\}_{i=1}^{h},$ $W_{1},W_{2},b_{1},b_{2}\}$, the computation of a transformer block is given by the combination of multi-head self-attention computation (Definition D.1) and feed-forward computation (Definition D.2). Formally, for an input matrix $X\in\mathbb{R}^{n\times d_{0}}$, we define:
@@
Definition D.4 (Reshape Layer).
@@
Definition D.5 (Complete transformer network).
We consider a transformer network as a composition of a transformer block (Definition D.3) with model weight $\theta^{h,m,r}$, which is:
@@
We especially say $\theta^{h,m,r}$ is the model weight that contains $h$ heads, $m$ hidden size for attention and $r$ hidden size for feed-forward. See Example D.6 for further explanation of the sequence-to-sequence mapping ${\cal F}$.
Example D.6.
We here give an example for the sequence-to-sequence mapping ${\cal F}$ in Definition D.5: Denote $L$ as the number of layers in some transformer network. For an input matrix $X\in\mathbb{R}^{n\times d}$, we use $E\in\mathbb{R}^{n\times d}$ to denote the positional encoding, we then define:
@@ D.2 Approximation via DiT
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
@@
Following Assumption B.10, we first denote $\widetilde{V}_{\tau}=\widetilde{{\cal M}}_{\tau}(c)$ for any $\tau\in[\frac{T}{\Delta t}]$ to discretize function ${\cal M}$. Then we have:
@@
where this step follows from Definition B.3 and Definition B.4.
@@
where these steps follow from Definition C.5 and simple algebras.
@@
where the first step follows the combination of Theorem C.10 and Eq. (D.2), and the differentiablity of $\widetilde{u}_{\tau}$ is ensure by Assumption B.7, the second step follows from Eq. (4).
@@
Appendix E Interpolation and Extrapolation
This section first introduce properties of HiPPO-LegS in Section E.1. Also, we bound the error of VLFM in Section E.2.
E.1 HiPPO-LegS Properties
Lemma E.1 (Proposition 6 in [30]).
@@ E.1 HiPPO-LegS Properties
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
  • Let Assumptions B.7, Assumption B.8, Assumption B.10 and Assumption B.9 hold.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
  • Let Assumptions B.7, Assumption B.8, Assumption B.10 and Assumption B.9 hold.
For any integer scale factor $\beta>0$, the frames of video $\widetilde{V}_{\tau}$ is scaled to $\widetilde{V}_{\beta\tau}$, it doesn’t affect the result of $H_{N}$ (Definition C.3).
@@
E.2 Error Bounds
Lemma E.3.
@@ E.2 Error Bounds
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ and matrix $G$ be defined as Definition C.4.
@@
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
  • Let Assumptions B.7, Assumption B.8, Assumption B.10 and Assumption B.9 hold.
@@
where the first step follows from Definition C.8, the second step follows from triangle inequality, the third step follows from $\sigma_{t}(\widetilde{u})\leq 1,\forall t\in[0,T]$ by some simple algebras and Definition C.6, the fourth step follows from the union bound of Gaussian tail bound (Fact B.11), the last step follows from the definition of $\epsilon_{2}$.
@@
where the first step follows from Definition C.5, the second step follows from optimal error of solving $\|MGH-Mu\|_{2}^{2}$, pesdueo-inverse matrix $(M\cdot G)^{\dagger}\in\mathbb{R}^{d\times\frac{T}{\Delta}}$ and defining a mask $M=\operatorname{diag}(m)$ where $m:=\{0,1\}^{\frac{T}{\Delta t}}$ and $\langle m,{\bf 1}_{\frac{T}{\Delta t}}\rangle=N$, the third step follows from the optimal error of solving $\|GH-u\|_{2}^{2}$, pesdueo-inverse matrix $G^{\dagger}\in\mathbb{R}^{d\times\frac{T}{\Delta}}$ and Lemma E.1, the fourth step follows from Cauchy–Schwarz inequality and the last step follows from the definition of $\epsilon_{2}$.
@@
where this step follows from Definition C.4, Fact B.12 and the definition of $\ell_{2}$ norm.
@@
where this step follows from Assumption B.9 and the definition of $\ell_{2}$ norm.
@@
where this step follows from Definition C.4 and the definition of $\ell_{2}$ norm.
@@
where the first step follows from Fact B.13, simple algebras, and Cauchy–Schwarz inequality, the second step follows from Eq. (6), Eq. (7), Definition C.4 and simeple algebras.
@@
these steps follow from simple algebras, Definition C.4 and the definition of $\epsilon_{3}$. ∎
Theorem E.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ and matrix $G$ be defined as Definition C.4.
@@
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition C.9.
  • Let Assumptions B.7, Assumption B.8, Assumption B.10 and Assumption B.9 hold.
@@
This proof follows from the combination of Assumption B.8, Theorem D.7 and Lemma E.3. ∎

### v2

@@ D.3 Training Objective
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
@@
Appendix E Diffusion Transformer
In this section, we first define the Diffusion Transformer in Appendix E.1. Moreover, we introduce the Approximation via DiT in Appendix E.2.
E.1 Definitions
Definition E.1 (Multi-head self-attention).
@@ E.1 Definitions
Definition E.2 (Feed-forward).
@@
Definition E.3 (Transformer block).
Given a set of model weights $\theta^{h,m,r}=\{\{(W_{Q}^{i},W_{K}^{i},W_{V}^{i},W_{O}^{i})\}_{i=1}^{h},$ $W_{1},W_{2},b_{1},b_{2}\}$, the computation of a transformer block is given by the combination of multi-head self-attention computation (Definition E.1) and feed-forward computation (Definition E.2). Formally, for an input matrix $X\in\mathbb{R}^{n\times d_{0}}$, we define:
@@
Definition E.4 (Reshape Layer).
@@
Definition E.5 (Complete transformer network).
We consider a transformer network as a composition of a transformer block (Definition E.3) with model weight $\theta^{h,m,r}$, which is:
@@
We especially say $\theta^{h,m,r}$ is the model weight that contains $h$ heads, $m$ hidden size for attention and $r$ hidden size for feed-forward. See Example E.6 for further explanation of the sequence-to-sequence mapping ${\cal F}$.
Example E.6.
We here give an example for the sequence-to-sequence mapping ${\cal F}$ in Definition E.5: Denote $L$ as the number of layers in some transformer network. For an input matrix $X\in\mathbb{R}^{n\times d}$, we use $E\in\mathbb{R}^{n\times d}$ to denote the positional encoding, we then define:
@@ E.2 Approximation via DiT
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
@@
Following Assumption C.10, we first denote $\widetilde{V}_{\tau}=\widetilde{{\cal M}}_{\tau}(c)$ for any $\tau\in[\frac{T}{\Delta t}]$ to discretize function ${\cal M}$. Then we have:
@@
where this step follows from Definition C.3 and Definition C.4.
@@
where these steps follow from Definition D.5 and simple algebras.
@@
where the first step follows the combination of Theorem D.10 and Eq. (E.2), and the differentiablity of $\widetilde{u}_{\tau}$ is ensure by Assumption C.7, the second step follows from Eq. (4).
@@
Appendix F Interpolation and Extrapolation
This section first introduce properties of HiPPO-LegS in Appendix F.1. Also, we bound the error of VLFM in Appendix F.2.
F.1 HiPPO-LegS Properties
Lemma F.1 (Proposition 6 in [33]).
@@ F.1 HiPPO-LegS Properties
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
  • Let Assumptions C.7, Assumption C.8, Assumption C.10 and Assumption C.9 hold.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
  • Let Assumptions C.7, Assumption C.8, Assumption C.10 and Assumption C.9 hold.
For any integer scale factor $\beta>0$, the frames of video $\widetilde{V}_{\tau}$ is scaled to $\widetilde{V}_{\beta\tau}$, it doesn’t affect the result of $H_{N}$ (Definition D.3).
@@
F.2 Error Bounds
Lemma F.3.
@@ F.2 Error Bounds
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ and matrix $G$ be defined as Definition D.4.
@@
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
  • Let Assumptions C.7, Assumption C.8, Assumption C.10 and Assumption C.9 hold.
@@
where the first step follows from Definition D.8, the second step follows from triangle inequality, the third step follows from $\sigma_{t}(\widetilde{u})\leq 1,\forall t\in[0,T]$ by some simple algebras and Definition D.6, the fourth step follows from the union bound of Gaussian tail bound (Fact C.11), the last step follows from the definition of $\epsilon_{2}$.
@@
where the first step follows from Definition D.5, the second step follows from optimal error of solving $\|MGH-Mu\|_{2}^{2}$, pesdueo-inverse matrix $(M\cdot G)^{\dagger}\in\mathbb{R}^{d\times\frac{T}{\Delta t}}$ and defining a mask $M=\operatorname{diag}(m)$ where $m:=\{0,1\}^{\frac{T}{\Delta t}}$ and $\langle m,{\bf 1}_{\frac{T}{\Delta t}}\rangle=N$, the third step follows from the optimal error of solving $\|GH-u\|_{2}^{2}$, pesdueo-inverse matrix $G^{\dagger}\in\mathbb{R}^{d\times\frac{T}{\Delta}}$ and Lemma F.1, the fourth step follows from Cauchy–Schwarz inequality and the last step follows from the definition of $\epsilon_{2}$.
@@
where this step follows from Definition D.4, Fact C.12 and the definition of $\ell_{2}$ norm.
@@
where this step follows from Assumption C.9 and the definition of $\ell_{2}$ norm.
@@
where this step follows from Definition D.4 and the definition of $\ell_{2}$ norm.
@@
where the first step follows from Fact C.13, simple algebras, and Cauchy–Schwarz inequality, the second step follows from Eq. (6), Eq. (7), Definition D.4 and simeple algebras.
@@
these steps follow from simple algebras, Definition D.4 and the definition of $\epsilon_{3}$. ∎
Theorem F.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ and matrix $G$ be defined as Definition D.4.
@@
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Let the training objective ${\cal L}(\theta)$ be defined as Definition D.9.
  • Let Assumptions C.7, Assumption C.8, Assumption C.10 and Assumption C.9 hold.
@@
This proof follows from the combination of Assumption C.8, Theorem E.7 and Lemma F.3. ∎